# A strong–weak duality for the 1d long-range Ising model

### v1

@@ Appendix B From continuum 1d LRI to AYK model
$u=\exp(-\rho^{2}(\theta/a+\kappa/\pi)/2)\;,\quad K=\rho^{2}/2\pi\;,\quad H/a=2\lambda_{1}\rho\;.$ (B.6)
Such relations are nonuniversal, so it is not surprising that in our derivation in the continuum they differ slightly from those obtained by Anderson and Yuval. However, they are qualitatively similar, and in particular they imply that the LRI model is a line in the $\{u,K\}$ plane, parametrized by $\rho^{2}$, because $\theta$ is an irrelevant coupling.
@@ C.3 Nonlinear sigma model formulation
In this formulation it is clear that the fundamental field is $\mathcal{U}(x)$, and not $\phi(x)$, which is introduced to solve the constraint and parametrize the circle. Notice that $\mathcal{M}(x)\equiv\bar{\mathcal{U}}(x)\partial_{x}\mathcal{U}(x)$ is the Maurer-Cartan form, so it is natural to express the kinetic term with it. In fact, if the theory was local, we would have a noninteracting action that is quadratic both in terms of $\mathcal{M}(x)$ and of $\mathcal{U}(x)$, because $\mathcal{M}(x)^{2}=-\partial_{x}\bar{\mathcal{U}}(x)\partial_{x}\mathcal{U}(x)$, by virtue of the constraint. In the long-range case instead it is quadratic only in $\mathcal{M}(x)$.
@@ Appendix D Correlators of $V_{\pm}$ and $\chi$ in the GFF
for generic $\delta$, where the normal ordering was defined in footnote 14, and we assume that $b_{0}\to 1$ for $\delta\to 0$.
@@ D.1 Insertions at infinity
Appendix E Logarithmic corrections to scaling near crossover
In this appendix, we study the logarithmic corrections to the scaling behavior of the critical 1d LRI Ising theory at the crossover. As discussed in the main text, at $s=1$ the flow to the IR fixed-point theory at $g=h=0$ is controlled by a marginally irrelevant operator, a linear combination of $\mathcal{O}_{g}$ and $\mathcal{O}_{h}$. A standard argument [110], also employed for the higher-dimensional LRI in appendix B of [9], shows that in the presence of marginally irrelevant operators, the CFT predictions for the IR behavior of correlators at criticality receive logarithmic corrections.
@@ F.1 One-point functions
$\langle\mathcal{O}_{g}(x)\rangle_{\mathcal{D}}$ $=\frac{g}{2}\int{\rm d}y\,\operatorname{tr}\langle\mathrm{P}\mathcal{O}_{g}(x)\mathcal{O}_{g}(y)\rangle_{0}=\frac{2g}{a}+O(g^{3},hg)\,.$
We can set the one-point function to zero by including a mixing term with the identity, i.e. we define the shifted operator $\mathcal{O}_{g}]$ as
@@ Appendix G The defect description of LRI
As demonstrated in [33], the $p$-dimensional LRI CFT can be realized as a co-dimension q=Dpq=D-p conformal defect for the free bulk massless scalar field $\Phi$, whihc propagates in the $D$-dimensional bulk. Reference [14] investigated the advantages of this construction to systematically constrain the dynamics of LRI using the defect conformal bootstrap approach [112].
@@ G.1 The defect description from the $\varphi^{4}$ formulation
$\Delta_{\varphi}=\frac{p-\epsilon}{4},\quad\Delta_{\varphi^{3}}=\frac{p+\epsilon}{4}\,$, (G.1)
where we have set $s=p/2+\epsilon$.
@@
$S=\int{\rm d}^{q}y\,{\rm d}^{p}x\frac{1}{2}(\partial\Phi)^{2}+\int{\rm d}^{d}x\frac{\lambda_{4}}{4}\varphi(x)^{4}\,$, (G.2)
with the condition that $\Phi(0,x)=\varphi(x)$, so that integrating out the $q$ transverse directions gives back the original non-local action. In this description, we have that $\Delta_{\Phi}=D/2-1=\frac{p-\epsilon}{4}$, and so the co-dimension of the defect is $q=1+\frac{p}{2}-\frac{\epsilon}{2}$.
@@
$\Delta_{+}=\frac{p-\epsilon}{4},\quad\Delta_{-}=\frac{p+\epsilon}{4}\,$, (G.4)
@@
It is interesting to note that the extension to the bulk necessarily comes with an ambiguity, corresponding to a change of boundary condition. For example, we are free to represent the same LRI as a defect in a different bulk theory, i.e. with the boundary condition $\Phi(0,x)=\varphi(x)^{3}$. Compared with the other boundary condition, this would be a dual conformal defect with co-dimension $q^{\prime}=4-q=3-\frac{p}{2}+\frac{\epsilon}{2}$ and $(\psi_{0}^{\pm})^{\prime}=\psi_{0}^{\mp}$.
@@ G.3 OPE relations
For certain special values of the parameters, the gamma functions in (G.3) develop poles and the OPE relations degenerate. For non-integer $q$ this only happens when
@@
Due to this poles, we must have that $\kappa_{1}(\Delta_{{T}},J)=\kappa_{2}(\Delta_{{T}},J)=0$ and hence $c_{11T}=c_{22T}=0$, while leaving $c_{12T}$ unconstrained. Hence, all odd-spin operators in $\psi_{1}\times\psi_{2}$ must have protected dimensions as above.

### v2

@@ Appendix B From continuum 1d LRI to AYK model
$\color[rgb]{0,0,0}g=\exp(-\rho^{2}(\theta/a+\kappa/\pi)/2)\;,\quad\mathcal{J}=\rho^{2}/2\pi\;,\quad H/a=2\lambda_{1}\rho\;.$} (B.6)
Such relations are nonuniversal, so it is not surprising that in our derivation in the continuum they differ slightly from those obtained by Anderson and Yuval. However, they are qualitatively similar, and in particular they imply that the LRI model without short-range term ($\theta=0$) and external field ($\lambda_{1}=0$) is a line in the $\{g,\mathcal{J}\}$ plane, parametrized by $\rho^{2}$.
@@ C.3 Nonlinear sigma model formulation
In this formulation the fundamental field is $\mathcal{U}(x)$, and not $\phi(x)$, which is introduced to solve the constraint and parametrize the circle. Notice that $\mathcal{M}(x)\equiv\bar{\mathcal{U}}(x)\partial_{x}\mathcal{U}(x)$ is the Maurer-Cartan form, so it is natural to express the kinetic term with it. In fact, if the theory was local, we would have a noninteracting action that is quadratic both in terms of $\mathcal{M}(x)$ and of $\mathcal{U}(x)$, because $\mathcal{M}(x)^{2}=-\partial_{x}\bar{\mathcal{U}}(x)\partial_{x}\mathcal{U}(x)$, by virtue of the constraint. In the long-range case instead it is quadratic only in $\mathcal{M}(x)$.
@@ Appendix D Correlators of $V_{\pm}$ and $\chi$ in the GFF
for generic $\delta$, where the normal ordering was defined in footnote 15, and we assume that $b_{0}\to 1$ for $\delta\to 0$.
@@ D.1 Insertions at infinity
Appendix E Logarithmic corrections to scaling at the crossover
In this appendix, we study the logarithmic corrections to the scaling behavior of the critical 1d LRI Ising theory at the crossover. As discussed in the main text, at $s=1$ the flow to the IR fixed-point theory at $g=h=0$ is controlled by a marginally irrelevant operator (a linear combination of $\mathcal{O}_{g}$ and $\mathcal{O}_{h}$). A standard argument [61], also employed for the higher-dimensional LRI in appendix B of [10], shows that in the presence of marginally irrelevant operators, the CFT predictions for the IR behavior of correlators at criticality receive logarithmic corrections.
@@ F.1 One-point functions
$\langle\mathcal{O}_{g}(x)\rangle_{\mathcal{D}}$ $=\frac{g}{2}\int{\rm d}y\,\operatorname{tr}\langle\mathrm{P}\mathcal{O}_{g}(x)\mathcal{O}_{g}(y)\rangle_{0}=\frac{2g}{a}+O(g^{3},hg)\,.$ (F.1)
We can as usual subtract the power-law divergence, and thus set the one-point function to zero by including a mixing term with the identity, i.e. we define the shifted operator $\mathcal{O}_{g}]$ as
@@ Appendix G The defect description of LRI
As demonstrated in [33], the $p$-dimensional LRI CFT can be realized as a co-dimension q=Dpq=D-p conformal defect for the free bulk massless scalar field $\Phi$, which propagates in the $D$-dimensional bulk. Reference [14] investigated the advantages of this construction to systematically constrain the dynamics of LRI using the defect conformal bootstrap approach [115].
@@ G.1 The defect description from the $\varphi^{4}$ formulation
$\Delta_{\varphi}=\frac{p-\epsilon}{4},\quad\Delta_{\varphi^{3}}={\color[rgb]{0,0,0}p-\Delta_{\varphi}}\,$, (G.1)
where we have set $s=(p+\epsilon)/2$.
@@
$S=\int{\rm d}^{q}y\,{\rm d}^{p}x\frac{1}{2}(\partial\Phi)^{2}+\int{\rm d}^{p}x\frac{\lambda_{4}}{4}\varphi(x)^{4}\,$, (G.2)
with the condition that $\Phi(0,x)=\varphi(x)$, so that integrating out the $q$ transverse directions gives back the original non-local action. In this description, we have that $\Delta_{\Phi}=D/2-1=\frac{p-\epsilon}{4}$, and so the co-dimension of the defect is $q={\color[rgb]{0,0,0}2-\frac{p}{2}-\frac{\epsilon}{2}}$.
@@
$\Delta_{+}=\frac{p-\epsilon}{4},\quad\Delta_{-}={\color[rgb]{0,0,0}p-\Delta_{\varphi}}\,$, (G.4)
@@
It is interesting to note that the extension to the bulk necessarily comes with an ambiguity, corresponding to a change of boundary condition. For example, we are free to represent the same LRI as a defect in a different bulk theory, i.e. with the boundary condition $\Phi(0,x)=\varphi(x)^{3}$. Compared with the other boundary condition, this would be a dual conformal defect with co-dimension $q^{\prime}=4-q$ and $(\psi_{0}^{\pm})^{\prime}=\psi_{0}^{\mp}$.
@@ G.3 OPE relations
By Bose symmetry, for $J$ odd we must have that $c_{11T}=c_{22T}=0$, while leaving $c_{12T}$ unconstrained. Combining with the OPE relations (G.3), for non-integer $q$ this condition requires that:
@@
Hence, all odd-spin operators in $\psi_{1}\times\psi_{2}$ must have protected dimensions as above.